OpenStudy (lilkg77):

Suppose that G(x) = F(x + 4) + 2. Which statement best compares the graph of G(x) with the graph of F(x)? A. The graph of G(x) is the graph of F(x) shifted 4 units to the right and 2 units down. B. The graph of G(x) is the graph of F(x) shifted 4 units to the left and 2 units up. C. The graph of G(x) is the graph of F(x) shifted 4 units to the right and 2 units up. D. The graph of G(x) is the graph of F(x) shifted 4 units to the left and 2 units down.

OpenStudy (agent0smith):

Any ideas?

OpenStudy (lilkg77):

Not sure

OpenStudy (agent0smith):

What does the +2 do to the graph?

OpenStudy (lilkg77):

It goes right @agent0smith

OpenStudy (mrhoola):

The graph should shift up by 2 units , not to the right

OpenStudy (mrhoola):

additionally , your answer is not listed in ,A,B,C,D

OpenStudy (mertsj):

Choice B is the correct answer.

OpenStudy (mathmate):

@lilkg77 g(x)=f(x-h)+k moves the graph of f(x) h units to the \(right\) (but note the minus sign of h) and k units up. for example, g(x)=f(x+3)-6 =f(x-(-3))+(-6), so h=-3, k=-6, moves the graph of f(x) 3 units to the left and 6 units down.